OpenStudy (anonymous):

the point (3,4) is the midpoint of (1, 10) and (5, k) find the value of k ??

OpenStudy (shamim):

u know the formula of middle point is x=(x1+x2)/2

OpenStudy (shamim):

nd y=(y1+y2)/2

OpenStudy (anonymous):

i do y1-y2 over x2-x1 ??

OpenStudy (anonymous):

how do i find k?

OpenStudy (shamim):

no

OpenStudy (shamim):

(5+1)/2=3 right

OpenStudy (anonymous):

yes

OpenStudy (shamim):

its for x

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes

OpenStudy (shamim):

need to give u an attachment

OpenStudy (anonymous):

ok thanks

OpenStudy (shamim):

feel free to ask if u hv any confusion